OpenStudy (anonymous):

2/3n-5/2=-3.... what is n?? plz help

OpenStudy (whpalmer4):

\[\frac{2}{3}n - \frac{5}{2} =-3\]Is that your equation?

OpenStudy (whpalmer4):

\(n\) is the letter between \(m\) and \(o\), of course :-)

OpenStudy (whpalmer4):

I would suggest that you multiply everything in the equation by 3*2=6. That will get rid of the fractions and leave you with an easy equation to solve for \(n\).

OpenStudy (anonymous):

add the 5/2 on the right side, than multiply by 3/2 to get n.

OpenStudy (anonymous):